OpenStudy (anonymous):

Find the value of y in the tessellation shown. A) 45° B) 90° C) 135° D) 180°

OpenStudy (pfenn1):

Are you going to show the tessellation?

OpenStudy (anonymous):

yea trying to attach it

OpenStudy (anonymous):

ok i sent it

OpenStudy (anonymous):

hello ?

OpenStudy (pfenn1):

If I understand the question correctly, it is asking for the interior angle of a regular octagon.

OpenStudy (anonymous):

yea that's what i don't understand it just says find y

OpenStudy (anonymous):

what did you get ?

OpenStudy (anonymous):

i got a but idk how i got that

OpenStudy (pfenn1):

The general formula for figuring the interior angle of a regular polygon is\[Interior \angle= \left( n-2 \over n \right)\times \left( 180 \over n \right)\]where n is the number of sides of the polygon. From http://www.mathsisfun.com/geometry/regular-polygons.html

OpenStudy (anonymous):

thany you i am soo ready for this EOCT bc of you thank you sooo much

OpenStudy (pfenn1):

you're welcome